OpenStudy (anonymous):

if sqrt(a-b) = (b/a)^(1-2x) then find the value of x

OpenStudy (anonymous):

Solve for x: sqrt(a-b) = (b/a)^(1-2 x) Reverse the equality in sqrt(a-b) = (b/a)^(1-2 x) in order to isolate x to the left hand side. sqrt(a-b) = (b/a)^(1-2 x) is equivalent to (b/a)^(1-2 x) = sqrt(a-b): (b/a)^(1-2 x) = sqrt(a-b) Taking reciporicals of both sides lets us solve for x in the numerator. Take reciporicals of both sides: (a (b/a)^(2 x))/b = 1/sqrt(a-b) Divide both sides by a constant to simplify the equation. Divide both sides by a/b: (b/a)^(2 x) = b/(a sqrt(a-b)) Eliminate the exponential from the left hand side. Take the logarithm base b/a of both sides: 2 x = (log(b/(a sqrt(a-b))))/(log(b/a))+((2 i) pi n)/(log(b/a)) for n element Z Solve for x. Divide both sides by 2: Answer: | | x = (log(b/(a sqrt(a-b))))/(2 log(b/a))+(i pi n)/(log(b/a)) for n element Z

OpenStudy (anonymous):

can it be shown in a simplified form so dat i can undrstnd? u might like to use "equation"(symbols) in order to clarify